NAOJ GW Elog Logbook 3.2

Also last week we found out the lens combination did not work well because I made some mistake in the calculation, so we tried to measure the better shape first order again, and found a new solution. The second lens should be put 40.1cm from the MZ beam splitter and the third 62cm. In this new configuration, all the lenses are in front of Faraday, so we don't need to worry about the affect to the reflection beam.
The other problem we mentioned in entry 457 is that all the optics on the 2nd FI path are not along the hole, so we can not screw the rail on the bench, so we used clamps. The path finished shows in the pictures, the way of clamping the rail now does not look very fixed, we will improve it later.
Since the new beam splitter has been installed, the green power injected into the cavity now is around 8mW.
With everything settling down, we could inject the beam into the chamber and align the cavity again. But today we are not able to lock the cavity, one reason is that the input mirror local control cannot hold the mirror fixed for long time, so while we were trying to find better mode matching, we continued to lose the alignment of the input. The other side, the beam produced by the new configuration is different from before, according to the calculation, the beam size arrive at 2 inch mirror should be close to our previous one, but the beam from the bench is much smaller than before and reflection of the input is larger than before. It could be the reason why we always have some Laguerre modes. So tomorrow we can try to change the driving matrix of the input and PR mirror, get better control of them and try to see if we are able to lock or not.
https://www.dropbox.com/s/nerrc4i4yx3bw8v/VID_1.MOV?dl=0

When we tried to study the MC finesse we did a mistake.
As we can see on the picture, it seems that the two TEM00 peaks are on the same part of the ramp (meaning same sign of cavity change length). But there was an offset as the TEM00 peaks show some small peaks symetric with respect to TEM00 peak.
So we took new data (second plot).
Finally we could extract the following parameters using the following equations : T0 / ( 1 + 4r/(1-r²) * sin ( PI (x - x0)/ FSR ) ² )
T0 = 0.98
FSR = 8.2072 ms
bg = 0.0205
r = 0.9902
x0 = -1.617 ms
with a R² factor = 0.9929
We could mesure the FMWH of the peaks at 0.026 ms which give us a finesse = 315.66

With all this reference today we started to install the AOM. The first lens has 100mm focal length which put as close as the original beam splitter. Then installed a new beam splitter after this lens and align both the transmission and reflection path of it.
The AOM was put after the mirror to get a good beam size, with the steering mirror, I aligned the first order straight first and then by adjusting the stage to increase the efficiency of the first order. It took long time for us to adjust the AOM, but still cannot get a perfect round beam, then we check the beam entering the AOM, from there the beam already has some strange shape halo around it. But before the beam splitter, we did not see this strange things, I think this is what we expected to see here, since when we installed the original third lens of the telescope, we already found out the shape affected a lot by the lens position. We will try to align better.
Just for testing the new telescope works or not, we put the other two lenses at their position, but it seems this combination does not works well as we tested it. One of the reason I think it may because when we did the test, the first lens position is not as close as now to the beam splitter, but I preferred to keep the first lens position like it is now, since this one has smaller divergence. This means we need to find other lenses, so we measured the beam size of the first order again after the AOM. Now the result I got is that we need to put the other two 250mm lens, one is at 40.5cm, the other at 88cm, which is doable.
Another problem we found it is that all the optics on the second Faraday path is not along the holes. So if we put the lens on the rail, the beam will be quite off-center and the range we can move with the screw is not enough to let the beam pass through the lens center.


At the beginning of this week we proceed some extra checks of local control problems.
First of all we get some transfer functions of some mirrors. Attached to this entry are some strange TF. For example the PR TF shows that when we inject noise in yaw, there is a correlation with pitch. But this correlation doesn't seem to appear when we inject noise in pitch. Also, IM TF shows correlation between pitch and yaw wherever we send noise.
This confirms our first idea that one magnet of PR has probably fallen.
We then proceed to do some others tests.
First we used the "normal vi" and not the latest version which is not able to get a proper feedback.
Second, we check what happend when sending noise to each coil :
The coil number 3 of PR doesn't respond.
All BS coils seem ok but every coil impact both pitch and yaw (mainly yaw)
IM coils (2,4) move only yaw , coil 3 move only pitch but coil 1 doesn't respond.
EM coils seem good : coils (1,3) move only yaw, coils (2,4) move only pitch
Third we manage to align roughly the cavity again using picomotors.
Finally, after removing offset of local control, we could align the cavity using local control.
We locked the cavity but the stability doesn't seem to last for very long.
Finally, we could get all the references needed to be able to align the beam with the AOM.

Yesterday (August 7th 2017) , in order to get references, we tried to lock the cavity. However, it was impossible to control some of the mirrors due to saturation of the correction signal.
Last lock was effected the previous monday ( July 31st 2017 ).
There were some earthquake in the mean time ( August 3d, 4th 2017) which we suspect removed one magnet from the PR suspended mirror at least ( if we excite one of the coil of this mirror there is no change in error signal from local control, 3 others coil shows some changes )
We also had some trouble with the BS suspended mirror but it seems to be due to an offset (-0.2 on yaw, 4.1 on pitch). Once removed, the results were coherent.
Down are listed the results of some of the test we did to chack the saturation problem :
open-loop (YAW;PITCH) | correction with open-loop offset value (YAW;PITCH) | max/min values for saturating the correction | |
PR | 0.242 ; -2.6 | 0 +/- 0.1 ; 0 +/- 1.5 | 2.945/-5 ; -1.2/3.9 (high correction values before + every changes saturates (peak) ) |
BS | -0.44 ; 0.54 | 0+/- 0.1 ; 0 +/- 0.5 | 1.72/-2.52 ; 1.065/-0.835 |
IM | -0.356 ; 0.145 | 0 +/- 0.01 ; 0+/- 0.2 | ok/ok ; ok/-1.14 |
EM | 0.55 ; 3.04 | 0.27 +/-0.1 ; 2.1 +/-0.1 | ok/ok ; ok/ok |
It also seems that the EM has some trouble ( pitch correction with open-loop offset is really different from 0 )
For both IM and EM there was "PSD tilt angle = -0.04 "

As usual, we tested the AOM on the transmission of the beam splitter. From our last test the AOM should be put 17.5cm from the beam splitter, but after checking, there is not enough space to put the AOM before the mirror, as close as we could to the mirror we can put the AOM 25cm from the beam splitter. And then if we put the second and third lens more or less at the same position as our last test, when the beam arrives at the position of FI, the diffraction orders are not separated enough to filter by the aperture. Then we tried to move the two lenses, situation did not get better.
So our idea is to filter the beam after the AOM but before the second lens. So we tried to do this and wanted to check the beam size after so we can find another proper combination of lens. When we measure the beam with the beam profiler, we found out the first order is very elliptical. We tried to align the AOM better, in the picture you can see there are some black lines which cut the beam into many pieces. Actually we had this kind of lines from the very beginning, but before we can get rid of them and get more or less a round first order with good alignment. This time we tired again and again, but still cannot get a good beam shape. We already sent an email to the company to ask if the black lines are only caused by bad alignment or something else. Meanwhile we checked the beam and found out the beam transmitted by the beam splitter itself has some tilt, and also the first 100mm lens could bring some astigmatism.
The other thing is that with the AOM position we decided today, we have only about 30cm to put lenses and recover the beam, although without the beam size measurement we cannot say if this space is enough or not, but if it is not, the other solution we considered is that after the FI and the waveplate, we have about 7.5cm space where we can put the lens. If we have to do this, one problem is that we can only take the reference outside the PR chamber, the other problem is that we are not sure if this will effect the beam reflect back to the PD which is used to lock the filter cavity.

After installed this mirror, the infrared became even weaker than before, I took a picture (pic 1)but it maybe too dim to see it. So when the cavity is not locked, we still can see the two infrared beam is flashing.( We are not sure why we can see two infrared beam very close to each other on the screen, maybe because some reflection between the two mirrors. ) When the cavity is locked the two beam both got stable. We tried to cut the infrared on the bench and we are sure what we saw is the infrared.
Then we tried to take a video of the infrared when the cavity is locked. This is the video: https://www.dropbox.com/s/gvsj0lsoei1pmqa/VID_1.MOV?dl=0
So this is the video I took when I asked Marc to jump in the central room, so you can see there are two beams moved to each other in the video.

To access to the MC finesse, we tried to use the following fit : T0/ ( 1 + 4R / (1-R)² * sin² ( pi (x-x0)/FSR ) ) + bg
Attached to this report are two results we could get : one without normalization, one with normalization.
By using Finesse = 4R / (1-R)², we could get the following results :
First case : 209
Second case (with normalization) : 360
By checking the ratio FSR over peak width : Finesse = 150.
We will try to undestand better how to fit this function

Today, it was not possible to lock the cavity because of a drift of the end mirror.
When we went to the end room to reset the local control, the temperature and humidity of the end room was quite high.
It appears that a air ventilation system was off. As soon as we turned it on, it started to feel more confortable.
After resetting the end mirror local control, it was possible to lock the cavity again.

I installed the 1310nm laser and the relative Imaging Unit
items:
- Laser controller
- fiber output
- golden half-inch mirror
- golden small prism mirror (before the cross point)
On the Imaging unit translation stage:
- golden large prism mirror (after the cross point)
- XY lens mount
- coated half ball
- Photo Detector

The Spectrum Analyzer provided data in ".DAT" file.
By using a program provided by Tatsumi-san, we were able to convert these data in ".DOT" file and then use Matlab to plot them.
This ".DOT" file is divided in 3 columns : frequency, magnitude and phase of the transfer function.
Attached to this entry are the open-loop, the electronic-loop and the optical loop.
They seem to be coherent with what the spectrum analyzer displayed during the measurement.

Before install everything, we checked the beam size after the transmission of the beam splitter, which is the same at the reflection. Take the origin at the front surface of the beam splitter, the beam waist is at -9.72cm, size is 53.36um. Then the first lens we use is 100mm, 5cm from the origin, but at the focal plane the beam diameter is larger than 1.3mm, which is the maximum the AOM can work according to the data sheet. So the AOM was put around 11cm, connected with the power, the AOM shows very clear diffraction orders.
Then another 100mm lens was put at 25cm. After putting this lens, we did some measurement of the beam size, and found the position to put the third lens in order to have a good size at the 2inch mirror of the telescope(less than 1.3mm). The third lens is 175mm, was put at 48.5cm. Then we checked the beam shape far, there is no obvious astigmatism as far as we can see. Measuring the beam again and we got the result that the beam at 2 inch mirror should be less than 1.2mm.
But then we found out if we want to change the MZ design and put another beam splitter after the one we have now, the first 100mm lens should be put further to give enough space to the BS. With Jammt, I did the simulation and found out if we move everything together 2.5cm further from the position mentioned before, the 175mm lens was not capable to focus the beam enough on the 2inch mirror. I tried other focal length, the 200mm should work.
The other problem is after moving(pic 1), the beam after the first lens will diverge more compared to the previous design(pic 2),sSince the two green dash lines in two pictures have the same distance. Divergence of the beam is one of the question we concern most, so I checked on the bench, after changing the AOM still works well.
It seems this new configuration is acceptable, we are going to install another beam splitter first and start to change the lenses.

The amplitude of the loop transfer functions plotted so far are actually the square of the real amplitude. The problem comes from the way I treated data saved by the spectrum analyzer. Each file is composed of 3 columns: frequency, real part (a) and imaginary part (b) of the TF. Of course amplitude and phase are recovered by doing:
Amplitude = sqrt (a^2 +b^2)
Phase = angle (a+i*b)
Due to an oversight, I had replaced the sqare root with the absolute value in the amplitude computation. This explain the unexpected behaviour (1/f^2 instead of 1/f) of the openloop TF around the UGF.
We will upload soon new TFs measurements (taken by Yuefan and Marc on monday night) properly plotted.

Mode cleaner cavity consists of three mirrors, two of them are flat and one is curve. After the curve mirror there is the PZT to change the cavity length for finding a good mode matching. This piezo is connected to the output of PZT driver whose input is connected to a function generator to provide the scan signal.
Since we only want to do a simple test, so we did not use the telescope design but only one 200mm lens after the beam splitter, then two steering mirrors used to align the cavity, at the output of the MC, a PD with DC output is used to see the modes. The whole configuration shows in pic 1.
We used 25Hz ramp wave with amplitude of 1Vpp to scan the cavity. At the beginning, we only saw some fluctuation but no peaks. When we tried to make the output beam go straight, we were not able to do it.(Always cut by the mirror mount) So we removed the MC and aligned better from the lens, sent the beam after the mirror far enough to make sure it goes along the holes of table. When we put back the MC, we could see some higher modes at the output and also the curve mirror has some transmission beam this time. Put back the PD, we saw pic 2 on the oscilloscope. By checking the beam shape with the curve mirror transmission and the spectrum, we got better mode matching. In pic 3, the highest peak is TEM00, we also checked it by moving the voltage of the PZT driver by hand. I think this means the mechanical part of the MC works well, we are able to align the cavity with this design.

A small problem was encounter : While screwing the cover part of the hole of cavity to the mode cleaner, a screw stayed stuck.
This was due to the fact that too long screw were proposed on the design. Instead of using M4*12, we used M4*8 screw for this part. We also used ethanol to avoid to stuck another screw.
Also, we couldn't find M3*16 screws so we used M3*15 screws.
To protect the wire used for the piezo power supply, we used a small piece of aluminum folded. One part is screwed to the optical bench while the other part hold a adaptor between BNC and the 2 wires for the piezo.

The table is to be updated with the values of yesterday (in blue).
low probe power | high probe power | |||
signal | AC | 23mV | AC | 230mV→295mV |
DC | 440mV | DC | 5400mV→5200mV | |
AC/DC | 0.052 | AC/DC | 0.043→0.057 | |
noise With sample |
AC_rms | 0.1mV | AC_rms | 2.8mV→1mV |
DC | 440mV | DC | 5400mV→5200mV | |
AC_rms/DC | 2.30E-04 | AC_rms/DC | 4.5e-4→1.9e-4 | |
ppm | 884ppm | ppm | 2000ppm→667ppm | |
noise Without sample |
AC_rms | 3μV | AC_rms | 70μV→ 200μV |
DC | 0.65V | DC | 6.5V | |
AC_rms/DC | 4.60E-06 | AC_rms/DC | 1e-5→3e-5 | |
ppm | 18ppm | ppm | 46ppm→108ppm |

This week, in order to check the AOM characteristics, we install the AOM after a beam splitter on the green path. By using a beam splitter before the AOM and 2 powermeters ( 1 one reflexion, the other on the transmission at the output of the AOM ) and checking their ratio, we were able to characterize the AOM despite still having power fluctuations on the green beam. The optical setup used is described in an attached figure.
By changing the RF power send to the AOM, we were able to characterize the AOM 1st order with the use of a gaussian fit ( even if this wasn’t really a gaussian, it helped to locate the maximum) as following :
- Maximum efficiency : 73 % @ RF Power 28.4 dBm ( 692 mW)
The AOM test sheet said that we could expect a 1st order efficiency superior than 85% at 633 nm. In this case, our alignment was approximative as we wanted to check only the response of the AOM to different RF power.
Then we tried to put the AOM on the right position on the optical bench. As the AOM need a small input beam size, we put it in the middle of 2 lenses ( f = 100 mm ) .
At that position, we couldn't see anymore any diffraction order.
First, we checked the green Power Density sent to the AOM. We measure 10W/mm² when the AOM test sheet limit this power density to 2.5 W/mm². Hopefully, we reduced quickly (after few minutes) the laser power down to 2 W/mm². In regard to this, we contact AA Opto-Electronic, manufacturer of this AOM. Following their advice, we check that the crystal was still transparent without any visible damages.
Then, we tried to put the AOM back on the characterization position. We were able to see again diffraction orders. We realize again the characterization of the 1st order efficiency and obtain :
- Maximum efficiency : 69 % @ RF Power 28.3 dBm ( 676 mW) We expect that the difference might be due to misalignment.
After that, we checked the polarization of the green beam using a PBS because this AOM needs a vertical polarization to work. We found that in both positions the green beam has a vertical polarization as we expect.
The last difference is the divergence of the beam. Indeed the beam is really more diverging in the right position (5.6 mrad) than on the characterization position (1.6 mrad) compared to the diffraction angle (16.6 mrad).
To correct this problem we will try to change the lenses configuration in order to obtain a smaller divergence of the beam on the right AOM position.

case | std(AC) | DC | std(AC/DC) | AC | std(AC) | AC/DC | std(AC/DC) | std(ppm) | std(ppm) |
pump | OFF | OFF/ON | OFF | ON | ON | ON | ON | OFF | ON |
without IR filter | 3.4uV | 0.77V | 4.4uV | 0.042V | 128uV | 0.0545 | 1.67E-04 | 16 | 621 |
with IR filter | 2.7uV | 0.45V | 6.1uV | 0.024V | 99.5uV | 0.053 | 2.26E-04 | 23 | 873 |
We found out that the DC changes from 0.46V to 0.44V when switching off the pump. This happens only when there is the sample, this means that some pump is scattered from the sample.
low probe power | high probe power | |||
signal | AC | 23mV | AC | 230mV |
DC | 440mV | DC | 5400mV | |
AC/DC | 0.052 | AC/DC | 0.043 | |
noise With sample |
AC_rms | 0.1mV | AC_rms | 2.8mV |
DC | 440mV | DC | 5400mV | |
AC_rms/DC | 2.30E-04 | AC_rms/DC | 4.50E-04 | |
ppm | 884ppm | ppm | 2000ppm | |
noise Without sample |
AC_rms | 3μV | AC_rms | 70μV |
DC | 0.65V | DC | 6.5V | |
AC_rms/DC | 4.60E-06 | AC_rms/DC | 1.00E-05 | |
ppm | 18ppm | ppm | 46ppm |
Today we checked again the signal level at the above conditions and we found almost the same values of the table above but the AC noise with low probe power and without sample was higher: around 200 μV instead of 70μV.
Then Kuroki suggested to cover the Imaging Unit to protect from wind (as it was in the original setup last year) and the noise became between 50-100μV , then we removed the cover and the noise remained on the same level 50-100μV. We think we should cover better the optical parts, in order to avoid temperature fluctuations which might affect the noise.
The table is to be updated with the values of yesterday (in blue).
low probe power | high probe power | |||
signal | AC | 23mV | AC | 230mV→295mV |
DC | 440mV | DC | 5400mV→5200mV | |
AC/DC | 0.052 | AC/DC | 0.043→0.057 | |
noise With sample |
AC_rms | 0.1mV | AC_rms | 2.8mV→1mV |
DC | 440mV | DC | 5400mV→5200mV | |
AC_rms/DC | 2.30E-04 | AC_rms/DC | 4.5e-4→1.9e-4 | |
ppm | 884ppm | ppm | 2000ppm→667ppm | |
noise Without sample |
AC_rms | 3μV | AC_rms | 70μV→ 200μV |
DC | 0.65V | DC | 6.5V | |
AC_rms/DC | 4.60E-06 | AC_rms/DC | 1e-5→3e-5 | |
ppm | 18ppm | ppm | 46ppm→108ppm |

We tried to calibrate the AOM again. Since the PD has too much effect on the power fluctuation, we decided to put two power meter in 0 and 1st order, so when we change the RF power, we can see the difference between these two orders. But we cannot find a good position to put the power meter that two order is separated enough and also the beam size is smaller than the aperture of the power meter. So we put another beam splitter after the MZ BS, and two power meter on two path of this beam splitter. Then with spectrum analyzer, we did the ratio between these two power in time domain to see the real change effected by the RF power.